Sir, please solve:
At a certain instant, a piece of radioactive material contains 1012 atoms. The half-life of the material is 30 days.
(i) Find the number of disintegrations in the first second.
(ii) What time would elapse before 104 atoms remain?
(iii) What is the count rate after this time?

Grade:12

2 Answers

Abhishek Kumar
9 years ago
half life = 1/\lambda
so, \lambda = 1/(30 X24 X 60 X 60)
So, 1) dN = \lambda x 1012 x1 (\lambdaNdt)
2) If N = 104 = 1012 e-\lambdat
Solve for t, rest things are known.
3) Count rate = \lambdaN = \lambda104
Fareena
3 years ago
1) lambda= 0.693/t(1/2)
     lambda = 0.693/30       t(1/2)=30 days (given)
     lambda = 0.0231/day
     -dN/dt= N lambda
                = 10^12 × 0.0231
                = 2.31 × 10^10 disintegration/day
                = 2.31×10^10 ÷ 24×3600
                = 2.7× 10^5 disintegration/second